OpenStudy (anonymous):

6. Which ordered pair is a solution of the equation y = 3x? (1 point) (–2, –9) (–8, –18) (–8, –3) (–10, –30

OpenStudy (anonymous):

Please HELP! 6. Which ordered pair is a solution of the equation y = 3x? (1 point) (–2, –9) (–8, –18) (–8, –3) (–10, –30

OpenStudy (anonymous):

can you help Shames?

OpenStudy (whpalmer4):

Well, the ordered points are interpreted as (x,y). Take the first choice, (-2,-9) -> x = -2, y = -9, and plug it into the equation \(y = 3x\). Does that result in a true statement? \[-9 = 3(-2)\]\[-9 = -6\]Oops. Well, that's not the right choice! Rinse, lather, repeat...

OpenStudy (anonymous):

I think its D

OpenStudy (anonymous):

whatcha think Shames?

OpenStudy (whpalmer4):

Let's check: (-10,-30) -> x = -10, y = -30 y = 3x -30 = 3(-10) -30 = -30 looks good!

OpenStudy (anonymous):

i gotta another ?

OpenStudy (anonymous):

5. Which equation has no solution? (1 point) 8 – (5v + 3) = 5v – 5 3m – 6 = 5m + 7 – m 3w + 4 – w = 5w – 2(w – 2 )7y + 9 = 7y – 6

OpenStudy (anonymous):

any ideas?

OpenStudy (whpalmer4):

\[8-(5v+3) = 5v-5\]\[8-5v-3 = 5v-5\]\[5-5v = 5v-5\]Anything look suspicious about that to you?

OpenStudy (anonymous):

im trying to figure out how i got these wrong cuz im takeing summmer school

OpenStudy (anonymous):

the fact there aint an answer

OpenStudy (whpalmer4):

well, we haven't solved it yet. there is an answer — I just wanted to know if that looked weird or not. \[5-5v=5v-5\]Add 5v to each side \[5-5v+5v=5v+5v-5\]Add 5 to each side\[5+5-5v+5v=5v+5v-5+5\]Collect like terms\[10=10v\]\[v=1\]So that equation is okay, it has a solution.

OpenStudy (anonymous):

oh ok

OpenStudy (whpalmer4):

\[3m – 6 = 5m + 7 – m\]Let's collect like terms again \[3m-6 = 4m +7\]let's subtract 3m from both sides\[3m-6-3m=4m+7-3m\]subtract 7 from both sides\[3m-6-3m-7=4m+7-3m-7\]collect like terms again \[-13=m\]That's got a solution, so it's okay.

OpenStudy (anonymous):

Why do some people just come and watch isnt that creepy? too you

OpenStudy (whpalmer4):

\[3w + 4 – w = 5w – 2(w – 2)\]first use distributive property on the parentheses \[3w+4-w=5w-2w-2(-2)\]\[3w+4-w=5w-2w+4\]collect like terms\[2w+4=3w+4\]subtract 4 from both sides\[2w=3w\]subtract 2w from both sides\[0=w\]That has a solution.

OpenStudy (anonymous):

so its the last one right?

OpenStudy (whpalmer4):

Finally, \[7y + 9 = 7y – 6\]subtract 7y from both sides \[9=-6\]WTF? No solution there!

OpenStudy (anonymous):

what about my other question?

OpenStudy (whpalmer4):

there's another case you can have, which is that you end up with something like 0=0. That means that there are infinitely many solutions.

OpenStudy (whpalmer4):

what, about watching? I can't say that I notice, I'm busy making sure I don't make any mistakes. I frequently watch others post answers to interesting questions, so that I might learn something myself (a novel way of doing or explaining the problem, or even all new material).

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

i have a few more questions though if you dont mind

OpenStudy (whpalmer4):

and if the poster is making a mistake, I'll point it out, or if the explanation appears to be unduly confusing the asker, I may stick around and attempt to explain it a different way in hopes of helping them understand.

OpenStudy (whpalmer4):

best to post new questions in new posts...

OpenStudy (whpalmer4):

for one thing, people tend to stop looking at questions after a medal has been awarded (thinking they probably won't get one for any work they do there), so if the person you are working with gets called away, or gets bored, or stuck, you may not get anyone else to help...

OpenStudy (anonymous):

understood i put it up for you

OpenStudy (whpalmer4):

tag me there, please ("@whpalmer4")